# Relativistic energy-momentum tensor distributions in a polarized nucleon

### v1

@@ Abstract
We study in detail the relativistic distributions of energy, longitudinal momentum, longitudinal energy flux, and longitudinal thrust inside nucleons based on the quantum phase-space formalism. Similar to recent studies on the electromagnetic current, we include the effects of the nucleon polarization and show that the latter are essential for understanding how the Breit frame distributions transform under a longitudinal Lorentz boost. We also explicitly demonstrate that, in the infinite-momentum frame, these distributions allow one to recover not only the “good” but also the “bad” components of the light-front energy-momentum tensor distributions.
@@ I Introduction
In this work, we extend the analysis of Lorcé et al. (2019) by including the nucleon polarization. The paper is organized as follows. In Sec. II, we remind the parametrization of the EMT matrix elements and some of its key properties. Then, in Sec. III, we review the definition of the relativistic spatial distributions within the quantum phase-space approach, and we analyze the distributions of energy, longitudinal momentum, longitudinal energy flux, and longitudinal thrust for both unpolarized and transversely polarized nucleons. Switching to the LF components, we show in Sec. IV that these distributions coincide in IMF limit with the LF ones. After a summary of our results, we added three appendices where we (A) discuss some details regarding the value of the dipole mass used for quark spin FF, (B) present an alternative parametrization of the EMT matrix elements directly in terms of the BF multipole FFs, and (C) show additional plots for the longitudinal momentum and longitudinal thrust distributions.
@@ II Matrix elements of the energy-momentum tensor
The total EMT FFs are obtained by summing over the quark and gluon contributions
@@
where $\mathcal{F}=A,B,D,\bar{C},S$. Since the symmetrized total EMT is not renormalized Collins et al. (1977), the total FFs factors $A$, $B$, $D$, and $\bar{C}$ are scheme and scale independent. The conservation of total linear and angular momenta imply in addition the following constraints Leader and Lorcé (2014); Teryaev (1999); Lowdon et al. (2017); Cotogno et al. (2019)
@@
Such a form is supported, e.g., by calculations in the chiral quark-soliton model Goeke et al. (2007); Won et al. (2024a, 2022, 2023, b). The parameters of the multipole Ansatz for $A$, $B$, $\bar{C}$, and $D$ FFs are given in Ref. Lorcé et al. (2019). Regarding the quark spin FF, we will use the dipole mass $\Lambda_{S_{q}}=1.00$ GeV and the normalization $S_{q}(0)=\frac{1}{2}g_{A}^{0}=0.17$. Changing the value of this dipole mass strongly affects the sign of the OAM FF, the nucleon spin polarization, and the dipole structure of the longitudinal OAM distribution in the BF, see Appendix A for a detailed discussion.
@@ III Distributions in elastic frame
A class of Lorentz frames, called the elastic frame (EF), has been introduced in Ref. Lorcé et al. (2018) to study the relativistic spatial distributions of total angular momentum (TAM), orbital angular momentum (OAM), and spin inside the nucleon. The formalism developed into a relativistic phase-space picture Lorcé (2018a) and was then used to study the spatial distributions of the EMT Lorcé et al. (2019), the electromagnetic current Lorcé (2020); Kim and Kim (2021); Lorcé and Wang (2022); Chen and Lorcé (2022); Kim et al. (2023); Chen and Lorcé (2023); Hong et al. (2023), and the axial-vector current Chen et al. (2024); Chen (2024), along with their dependence on the hadron momentum.
@@ III.1 General expressions
$d\gamma=\frac{P^{0}}{\sqrt{P^{2}}},\qquad\beta=\frac{P_{z}}{P^{0}}.$ (16)
@@
with $E_{P}=\sqrt{P_{z}^{2}+M^{2}}$, and coincide with the forward limit $\Delta\to 0$ of the expressions in Eq. (16). It is clear from Eq. (17) that the spatial distributions $T^{00}$, $T^{03}$, $T^{30}$, and $T^{33}$ diverge in the IMF. Since our goal is to discuss the distortions induced by a Lorentz boost along the $z$-direction, we factor out the trivial Lorentz factors and define EMT distributions with $P_{z}$-independent normalizations. Namely, we consider similarly to Ref. Lorcé et al. (2019)
@@
The normalized distributions of energy $\rho_{a}$, longitudinal momentum $\mathcal{P}_{a}^{z}$, longitudinal energy flux $\mathcal{I}_{a}^{z}$, and longitudinal thrust $\Pi^{zz}_{a}$ can therefore be expressed as
@@ III.3 Longitudinal momentum and longitudinal energy flux distributions
where $L(Q^{2})=J(Q^{2})-S(Q^{2})$ is the OAM FF. Since the 3D BF momentum distributions
@@
In other words, the dipole moment in the BF momentum distribution is a direct manifestation of the internal OAM.
@@
III.4 Longitudinal thrust distribution
Total EF distribution of longitudinal thrust for a transversely polarized nucleon
Figure 5: Total EF frame distributions of longitudinal thrust in the transverse plane of a nucleon polarized along the $x$-axis for different values of the nucleon momentum. The total distributions are normalized to $M\beta^{2}_{P}$. Based on the simple multipole model of Ref. Lorcé et al. (2019) for the EMT FFs.
The spin-independent and spin-dependent amplitudes defining the longitudinal thrust distribution in the EF are given by
@@ III.4 Longitudinal thrust distribution
and the total EF distribution of longitudinal thrust is normalized as
@@
In the transverse BF, the longitudinal thrust amplitudes reduce to
@@
In Fig. 5, we show the total EF distributions of longitudinal thrust for non-zero $P_{z}$. The differences with the total EF energy distributions displayed in Fig. 3 are simply due to the different kinematical weighting of the contributions associated with $E_{a}$ and $F_{a}$, see Eqs. (30) and (52). This explains also why the dipole moment of the total longitudinal thrust distribution in the EF
@@ IV Distributions on the light front
Similarly to the analysis presented in Ref. Chen and Lorcé (2022) for the electromagnetic current, we can relate these LF amplitudes to the corresponding IMF amplitudes. Indeed, combining the results in Eq. (20), we find that
@@ V Summary
In this paper, we have studied the relativistic distributions of the local gauge-invariant and asymmetric energy-momentum tensor (EMT) current within the framework of the quantum phase-space formalism for the nucleons, taking into account for the first time the nucleon transverse polarization. We have focused on the “longitudinal” components of the EMT and defined the corresponding scaled distributions of energy, longitudinal momentum, longitudinal energy flux, and longitudinal thrust, including their decomposition into quark and gluon contributions. The frame dependence of these distributions has then been studied in detail and illustrated using a simple multipole model for the EMT form factors.
@@ Appendix A Quark axial singlet dipole mass
$\mathcal{P}_{a}^{z,P}\left(\bm{b}_{\perp},0\right)$ $=M\int\frac{d^{2}\Delta_{\perp}}{\left(2\pi\right)^{2}}\,e^{-i\bm{\Delta}_{\perp}\cdot\bm{b}_{\perp}}\frac{i\Delta_{y}}{2M}L_{a}(Q^{2}),$ (68)
where the OAM FF is defined by $L_{a}=J_{a}-S_{a}$. Since the gluon spin FF is zero, we consider only the quark part.
@@ Appendix B Breit frame multipole form factors
where $g^{\mu\nu}_{P}=g^{\mu\nu}-P^{\mu}P^{\nu}/P^{2}$ is the projector onto the subspace orthogonal to $P$. The force FF driving the longitudinal thrust in the transverse BF (54) is given by a linear combination of a monopole and a quadrupole FF, viz. $F_{a}=F_{0,a}-\frac{\tau}{3}F_{2,a}$ with $F_{0,a}=-\bar{C}_{a}-\frac{2}{3}\tau D_{a}$ and $F_{2,a}=D_{a}$.
Appendix C Spatial distributions of longitudinal momentum and longitudinal thrust
In Figs. 8 and 9, we show the quark and gluon contributions to the EF distributions of, respectively, longitudinal momentum and longitudinal thrust in an unpolarized nucleon at various values of $P_{z}$. Since these distributions are axially symmetric, only the radial profiles are displayed.
@@ Appendix C Spatial distributions of longitudinal momentum and longitudinal thrust
EF longitudinal thrust distributions for the unpolarized nucleon
@@
Figure 9: Radial EF distributions of longitudinal thrust in an unpolarized nucleon for different values of the nucleon momentum. The total longitudinal thrust distribution is normalized to $M\beta^{2}_{P}$. The left (right) column corresponds to the quark (gluon) contribution. The first row depicts the sum of the three contributions shown in the other rows. Based on the simple multipole model of Ref. Lorcé et al. (2019) for the EMT FFs.
Note that the longitudinal momentum distributions are purely induced by the Lorentz boost. In the absence of Wigner rotation, only the contributions associated with $E_{a}$ and $F_{a}$ would play a role. At large $P_{z}$, the contribution associated with $S_{a}$ is the only one that is suppressed kinematically. The longitudinal thrust distributions are similar to the energy distributions. The difference in the $P_{z}$-evolution between the contributions associated with $E_{a}$ and $F_{a}$ is essentially a reflection of the relative kinematical weight $\beta^{2}$ in Eqs. (20b) and (20f).
@@
we illustrate the same distributions for the nucleon polarized along the $x$-direction. Similarly to the polarized EF energy distributions, the dipole shift induced by the contribution associated with $J_{a}$ goes in the opposite direction to those associated with $E_{a}$ and $F_{a}$.
EF longitudinal thrust distributions for the transversely polarized nucleon
@@
Figure 11: EF distributions of longitudinal thrust at $b_{x}=0$ in a nucleon polarized along the $x$-axis for different values of the nucleon momentum. The total longitudinal thrust distribution is normalized to $M\beta^{2}_{P}$. The left (right) column corresponds to the quark (gluon) contribution. The first row depicts the sum of the three contributions shown in the other rows. Based on the simple multipole model of Ref. Lorcé et al. (2019) for the EMT FFs.

### v2

@@ Abstract
We study in detail the relativistic distributions of energy, longitudinal momentum, longitudinal energy flux, and axial momentum flux inside nucleons based on the quantum phase-space formalism. Similar to recent studies on the electromagnetic current, we include the effects of the nucleon polarization and show that the latter are essential for understanding how the Breit frame distributions transform under a longitudinal Lorentz boost. We also explicitly demonstrate that, in the infinite-momentum frame, these distributions allow one to recover not only the “good” but also the “bad” components of the light-front energy-momentum tensor distributions.
@@ I Introduction
In this work, we extend the analysis of Lorcé et al. (2019) by including the nucleon polarization. The paper is organized as follows. In Sec. II, we remind the parametrization of the EMT matrix elements and some of its key properties. Then, in Sec. III, we review the definition of the relativistic spatial distributions within the quantum phase-space approach, and we analyze the distributions of energy, longitudinal momentum, longitudinal energy flux, and axial momentum flux for both unpolarized and transversely polarized nucleons. Switching to the LF components, we show in Sec. IV that these distributions coincide in IMF limit with the LF ones. After a summary of our results, we added three appendices where we (A) discuss some details regarding the value of the dipole mass used for quark spin FF, (B) present an alternative parametrization of the EMT matrix elements directly in terms of the BF multipole FFs, and (C) show additional plots for the longitudinal momentum and axial momentum flux distributions.
@@ II Matrix elements of the energy-momentum tensor
According to General relativity (GR), the EMT is a symmetric rank-2 tensor. As a result, many early studies on the EMT in particle physics adopted the symmetric form, also known as the Belinfante-improved EMT Belinfante (1939); Pauli and Belinfante (1940). It should, however, be kept in mind that GR is a classical field theory, whereas the concept of spin, which is directly related to the antisymmetric part of the EMT, is of a quantum nature. Hence, restricting the analysis to the symmetric EMT alone is not well justified in the context of particle physics, see Ref. Leader and Lorcé (2014) for further details and discussions. In the present work, we consider the (more general) asymmetric form of the EMT, which allow us to decompose the total angular momentum (TAM) into spin and orbital angular momentum (OAM) contributions.
@@
The latter follows from the impossibility of writing down a local and gauge-invariant gluon spin operator Leader and Lorcé (2014). Consequently, the gluon TAM appears in a pure orbital form, which can be regarded as an “effective” one. The total EMT FFs are obtained by summing over the quark and gluon contributions
@@
where $\mathcal{F}=A,B,D,\bar{C},S$. Since the symmetrized total EMT is not renormalized Collins et al. (1977), the total FFs $A$, $B$, $D$, and $\bar{C}$ are scheme and scale independent. The conservation of total linear and angular momenta imply in addition the following constraints Leader and Lorcé (2014); Teryaev (1999); Lowdon et al. (2017); Cotogno et al. (2019)
@@
Such a form is supported, e.g., by calculations in the chiral quark-soliton model Goeke et al. (2007); Won et al. (2024a, 2022, 2023, b). The parameters of the multipole Ansatz for $A$, $B$, $\bar{C}$, and $D$ FFs are given in Ref. Lorcé et al. (2019). Regarding the quark spin FF, we will use the dipole mass $\Lambda_{S_{q}}=1.00$ GeV and the normalization $S_{q}(0)=\frac{1}{2}g_{A}^{0}=0.17$. Changing the value of this dipole mass strongly affects the sign of the OAM FF, $L_{a}(Q^{2})=\frac{1}{2}[A_{a}(Q^{2})+B_{a}(Q^{2})]-S_{a}(Q^{2})$, the nucleon spin polarization, and the dipole structure of the longitudinal OAM distribution in the BF, see Appendix A for a detailed discussion.
@@ III Distributions in elastic frame
A class of Lorentz frames, called the elastic frame (EF), has been introduced in Ref. Lorcé et al. (2018) to study the relativistic spatial distributions of TAM, OAM, and spin inside the nucleon. The formalism developed into a relativistic phase-space picture Lorcé (2018a) and was then used to study the spatial distributions of the EMT Lorcé et al. (2019), the electromagnetic current Lorcé (2020); Kim and Kim (2021); Lorcé and Wang (2022); Chen and Lorcé (2022); Kim et al. (2023); Chen and Lorcé (2023); Hong et al. (2023), and the axial-vector current Chen et al. (2024); Chen (2024), along with their dependence on the hadron momentum.
@@ III.1 General expressions
$\gamma=\frac{P^{0}}{\sqrt{P^{2}}},\qquad\beta=\frac{P_{z}}{P^{0}}.$ (16)
@@
with $E_{P}=\sqrt{P_{z}^{2}+M^{2}}$, and coincide with the forward limit $\Delta\to 0$ of the expressions in Eq. (16). It is clear from Eq. (17) that the spatial distributions $T^{00}$, $T^{03}$, $T^{30}$, and $T^{33}$ diverge in the IMF. Since our goal is to discuss the distortions induced by a Lorentz boost along the $z$-direction, we factor out the trivial Lorentz factors and define EMT distributions with $P_{z}$-independent normalizations. Namely, we consider similar to Ref. Lorcé et al. (2019)
@@
The normalized distributions of energy $\rho_{a}$, longitudinal momentum $\mathcal{P}_{a}^{z}$, longitudinal energy flux $\mathcal{I}_{a}^{z}$, and axial momentum flux (i.e. longitudinal flux of longitudinal momentum) $\Pi^{zz}_{a}$ can therefore be expressed as
@@ III.3 Longitudinal momentum and longitudinal energy flux distributions
where $L_{a}(Q^{2})=J_{a}(Q^{2})-S_{a}(Q^{2})$. Since the 3D BF momentum distributions
@@
In other words, the dipole moment in the BF momentum distribution is a direct manifestation of the internal OAM (recall that, in this work, the gluon spin contribution is effectively represented in an orbital form).
@@
III.4 Axial momentum flux distribution
Total EF distribution of axial momentum flux for a transversely polarized nucleon
Figure 5: Total EF frame distributions of axial momentum flux in the transverse plane of a nucleon polarized along the $x$-axis for different values of the nucleon momentum. The total distributions are normalized to $M\beta^{2}_{P}$. Based on the simple multipole model of Ref. Lorcé et al. (2019) for the EMT FFs.
The spin-independent and spin-dependent amplitudes defining the axial momentum flux distribution in the EF are given by
@@ III.4 Axial momentum flux distribution
and the total EF distribution of axial momentum flux is normalized as
@@
In the transverse BF, the axial momentum flux amplitudes reduce to
@@
In Fig. 5, we show the total EF distributions of axial momentum flux for non-zero $P_{z}$. The differences with the total EF energy distributions displayed in Fig. 3 are simply due to the different kinematical weighting of the contributions associated with $E_{a}$ and $F_{a}$, see Eqs. (30) and (52). This explains also why the dipole moment of the total axial momentum flux distribution in the EF
@@ IV Distributions on the light front
Similar to the analysis presented in Ref. Chen and Lorcé (2022) for the electromagnetic current, we can relate these LF amplitudes to the corresponding IMF amplitudes. Indeed, combining the results in Eq. (20), we find that
@@ V Summary
In this paper, we have studied the relativistic distributions of the local gauge-invariant and asymmetric energy-momentum tensor (EMT) current within the framework of the quantum phase-space formalism for the nucleons, taking into account for the first time the nucleon transverse polarization. We have focused on the “longitudinal” components of the EMT and defined the corresponding scaled distributions of energy, longitudinal momentum, longitudinal energy flux, and axial momentum flux, including their decomposition into quark and gluon contributions. The frame dependence of these distributions has then been studied in detail and illustrated using a simple multipole model for the EMT form factors.
@@ Appendix A Quark axial singlet dipole mass
$\mathcal{P}_{a}^{z,P}\left(\bm{b}_{\perp},0\right)$ $=M\int\frac{d^{2}\Delta_{\perp}}{\left(2\pi\right)^{2}}\,e^{-i\bm{\Delta}_{\perp}\cdot\bm{b}_{\perp}}\frac{i\Delta_{y}}{2M}L_{a}(Q^{2}).$ (68)
Since the gluon spin FF is zero, we consider only the quark part.
@@ Appendix B Breit frame multipole form factors
where $g^{\mu\nu}_{P}=g^{\mu\nu}-P^{\mu}P^{\nu}/P^{2}$ is the projector onto the subspace orthogonal to $P$. The force FF driving the axial momentum flux in the transverse BF (54) is given by a linear combination of a monopole and a quadrupole FF, viz. $F_{a}=F_{0,a}-\frac{\tau}{3}F_{2,a}$ with $F_{0,a}=-\bar{C}_{a}-\frac{2}{3}\tau D_{a}$ and $F_{2,a}=D_{a}$.
Appendix C Spatial distributions of longitudinal momentum and axial momentum flux
In Figs. 8 and 9, we show the quark and gluon contributions to the EF distributions of, respectively, longitudinal momentum and axial momentum flux in an unpolarized nucleon at various values of $P_{z}$. Since these distributions are axially symmetric, only the radial profiles are displayed.
@@ Appendix C Spatial distributions of longitudinal momentum and axial momentum flux
EF axial momentum flux distributions for the unpolarized nucleon
@@
Figure 9: Radial EF distributions of axial momentum flux in an unpolarized nucleon for different values of the nucleon momentum. The total axial momentum flux distribution is normalized to $M\beta^{2}_{P}$. The left (right) column corresponds to the quark (gluon) contribution. The first row depicts the sum of the three contributions shown in the other rows. Based on the simple multipole model of Ref. Lorcé et al. (2019) for the EMT FFs.
Note that the longitudinal momentum distributions are purely induced by the Lorentz boost. In the absence of Wigner rotation, only the contributions associated with $E_{a}$ and $F_{a}$ would play a role. At large $P_{z}$, the contribution associated with $S_{a}$ is the only one that is suppressed kinematically. The axial momentum flux distributions are similar to the energy distributions. The difference in the $P_{z}$-evolution between the contributions associated with $E_{a}$ and $F_{a}$ is essentially a reflection of the relative kinematical weight $\beta^{2}$ in Eqs. (20b) and (20f).
@@
we illustrate the same distributions for the nucleon polarized along the $x$-direction. Similar to the polarized EF energy distributions, the dipole shift induced by the contribution associated with $J_{a}$ goes in the opposite direction to those associated with $E_{a}$ and $F_{a}$.
EF axial momentum flux distributions for the transversely polarized nucleon
@@
Figure 11: EF distributions of axial momentum flux at $b_{x}=0$ in a nucleon polarized along the $x$-axis for different values of the nucleon momentum. The total axial momentum flux distribution is normalized to $M\beta^{2}_{P}$. The left (right) column corresponds to the quark (gluon) contribution. The first row depicts the sum of the three contributions shown in the other rows. Based on the simple multipole model of Ref. Lorcé et al. (2019) for the EMT FFs.